The music teacher is three times as old as Walt. In 12 years, she will be twice as old as Walt. How old is Walt?
Answer by mananth(16946)

The music teacher is three times as old as Walt. In 12 years, she will be twice as old as Walt. How old is Walt?
..
Let Walt be x years
teacher = 3x
...
after 12 years
Walt = x+12
teacher = 3x+12
..
3x+12=2(x+12)
3x+12=2x+24
-2x
3x-2x+12=2x-2x+24
x+12=24
-12
x=24-12
x=12 Walt's age.
teacher's age = 3x = 12*3=36 years
...
m.ananth@hotmail.ca